OpenStudy (anonymous):

I have a practice assignment before my big unit test tomorrow. I understand every subject except for volumes of a solid of revolution. I am going to post the questions from that section and if you could please show me the answers and how to get to them. I would greatly appreciate it! I really need to understand this concept! *In Problems 1-5, let R be the region bounded by y= (square root)x , y = (square root)1, and x =(square root)4.

OpenStudy (anonymous):

Each problem will describe a solid generated by rotating R about an axis. Write an integral expression that can be used to find the volume of the solid (do not evaluate). 1. The solid is generated by rotating R about the x-axis. 2.The solid is generated by rotating R about the y-axis. 3.The solid is generated by rotating R about y = 1. 4. The solid is generated by rotating R about x = −2. 5. The solid is generated by rotating R about y = 3. In Problems 1-5, let R be the region bounded by y = x ^2 + 1 and y = x + 1.Each problem will describe a solid generated by rotating R about an axis. Write an integral expression that can be used to find the volume of the solid (do not evaluate). 1. The solid is generated by rotating R about the y-axis. 2. The solid is generated by rotating R about the x-axis. 3. The solid is generated by rotating R about x = 1. 4. The solid is generated by rotating R about y = 2. 5. The solid is generated by rotating R about x = -3.

ganeshie8 (ganeshie8):

could u take a snapshot and attach if possible ?

ganeshie8 (ganeshie8):

*In Problems 1-5, let R be the region bounded by y= (square root)x , y = (square root)1, and x =(square root)4.

ganeshie8 (ganeshie8):

whats with : y = (square root)1, and x =(square root)4. ?

OpenStudy (anonymous):

I didn't know how to put in the square root button

OpenStudy (anonymous):

ill try to take picture hold on, its work problem in my online textbook. let me figure this out

ganeshie8 (ganeshie8):

take a screenshot of problem and attach the pic...

ganeshie8 (ganeshie8):

okay take ur time :)

OpenStudy (anonymous):

do you know how to take a screenshot because I don't...

ganeshie8 (ganeshie8):

install this : http://app.prntscr.com/

OpenStudy (anonymous):

are you still able to help me?

OpenStudy (anonymous):

@primeralph

ganeshie8 (ganeshie8):

start by sketching the region

ganeshie8 (ganeshie8):